OpenStudy (anonymous):

What is the y-intercept of the function y = 16x? A. 16 B. 1 C. y D. x

OpenStudy (anonymous):

For finding y intercept, put x = 0..

OpenStudy (anonymous):

i agree but that is not in the options

OpenStudy (anonymous):

Are you sure about your answer choices??

OpenStudy (anonymous):

Pretty sure its 0....

OpenStudy (anonymous):

Yeah y-intercept is zero because x is zero here..

OpenStudy (anonymous):

The answer is 1 ! So B

OpenStudy (anonymous):

Is that: \[\large y = 16^x\]

OpenStudy (anonymous):

If yes, then answer will be surely 1.. \[y = 16^0 = 1\]